# Training-Free Diffusion Priors for Text-to-Image Generation via Optimization-based Visual Inversion

### v1

@@ Abstract
Diffusion models have established the state-of-the-art in text-to-image generation, but their performance often relies on a diffusion prior network to translate text embeddings into the visual manifold for easier decoding. These priors are computationally expensive and require extensive training on massive datasets. In this work, we challenge the necessity of a trained prior at all by employing Optimization-based Visual Inversion (OVI), a training-free and data-free alternative, to replace the need for a prior. OVI initializes a latent visual representation from random pseudo-tokens and iteratively optimizes it to maximize the cosine similarity with input textual prompt embedding. We further propose two novel constraints, a Mahalanobis-based and a Nearest-Neighbor loss, to regularize the OVI optimization process toward the distribution of realistic images. Our experiments, conducted on Kandinsky 2.2, show that OVI can serve as an alternative to traditional priors. More importantly, our analysis reveals a critical flaw in current evaluation benchmarks like T2I-CompBench++, where simply using the text embedding as a prior achieves surprisingly high scores, despite lower perceptual quality. Our constrained OVI methods improve visual fidelity over this baseline, with the Nearest-Neighbor approach proving particularly effective, achieving quantitative scores comparable to or higher than the state-of-the-art data-efficient prior, indicating that the idea merits further investigation. The code will be publicly available upon acceptance.
@@ 1 Introduction
Text-to-image (T2I) generation has seen remarkable progress, with diffusion models StableDiffusion representing the current state-of-the-art. Architectures like unCLIP unCLIP have demonstrated that a hierarchical approach, which first generates a CLIP image embedding from a text caption using a prior model and then generates an image from that embedding using a decoder, can produce high-fidelity and diverse images. However, this performance comes at a significant cost: the diffusion prior is typically a large model that is expensive to train, requiring billions of image-text pairs and substantial computational resources.
@@
Recent work has focused on making these priors more efficient. ECLIPSE Eclipse, for example, introduced a data-efficient prior that achieves competitive results at a fraction of the parameters and training samples. Despite these advancements, the fundamental paradigm remains unchanged, the prior must still be learned through an extensive training phase, limiting accessibility and adaptability.
@@
In this work, we investigate a radical alternative: “can we entirely eliminate the training phase for the prior?” Building on recent work on modality inversion CrossTheGap, we leverage Optimization-based Visual Inversion (OVI), a data free and training-free method, originally introduced to cast intra-modal tasks into inter-modal ones, to substitute the prior with an on-the-fly optimization process. Instead of relying on a mapping network, OVI generates a CLIP image embedding by iteratively refining a random latent vector to maximize its semantic alignment with the target textual prompt embedding. To further guide this process towards plausible visual representations, we introduce two novel regularization strategies: a Mahalanobis constraint that pulls the embedding towards the global distribution of real image embeddings, and a Nearest-Neighbor constraint that aligns it with the most similar real image embedding from a reference dataset.
@@
Our preliminary results reveal several interesting aspects of T2I generation. We find that unconstrained OVI produces qualitatively similar results to directly using the text embedding in the decoder, a surprisingly strong baseline first noted in unCLIP. This baseline achieves state-of-the-art scores on the benchmark T2I-CompBench++ T2ICompBench, often exceeding an expensive trained priors like ECLIPSE. This suggests a potential “metric problem,” where current benchmarks may over-reward semantic similarity at the expense of visual quality and realism. Our constrained OVI methods, particularly the Nearest-Neighbor approach, improve visual fidelity and produce images almost competitive with ECLIPSE, suggesting that training-free optimization-based priors have merit and warrant further investigation as an alternative to their trained counterparts.
@@ 2.1 Diffusion Models for Text-to-Image Generation
Diffusion models have become the dominant architecture for T2I synthesis. The unCLIP unCLIP framework popularized a two-stage pipeline consisting of a prior and a decoder. The decoder, often a U-Net based model, is trained to generate an image conditioned on a CLIP image embedding. Our work leverages the publicly available Kandinsky 2.2 decoder Kandinsky, which is known for its high-quality image generation capabilities.
@@ 2.2 Image Priors for Text-to-Image Generation
The prior model is responsible for mapping a the textual prompt embedding $z_{y}$ to the corresponding image embedding $z_{x}$. The original unCLIP unCLIP explored both autoregressive and diffusion-based priors, with the latter generally proving more efficient and effective. More recently, ECLIPSE Eclipse proposed a lightweight, non-diffusion prior trained with a combination of a projection and a contrastive loss, significantly reducing computational requirements while maintaining high performance. Our work investigates the possibility of omitting the prior training phase, proposing a zero-shot, optimization-based substitute.
@@ 2.3 Modality Inversion and Optimization-based Methods
The idea of inverting a model or optimizing in a latent space is not new. In the context of CLIP, CrossTheGap explored modality inversion – Optimization-based Textaul Inverison (OTI) and Optimization-based Visual Inversion (OVI) – to expose and bridge the gap between the text and image embedding manifolds. In parallel, methods like SEARLE SEARLE have used optimization to project a reference image into a “pseudo-token” for composed image retrieval, demonstrating the power of training-free optimization to achieve strong multimodal alignment. Our OVI-based approach builds on these concepts, adapting modality inversion to serve as a generative, training- and data-free prior for text-to-image generation.
@@ 3.2 Optimization-based Visual Inversion (OVI)
The optimization-based approach frames the prior’s task as an optimization problem. Instead of learning a mapping, we directly search for an image embedding $\hat{z}_{x}$ that aligns with the given text embedding $z_{y}$. We initialize a set of NN learnable pseudo-tokens, which form our latent image embedding $z_{img}^{(0)}$. We then iteratively update this embedding for $T$ steps using an optimizer to minimize a loss function based on cosine similarity with the target text embedding, as computed by the CLIP encoders:
@@ 3.3 Mahalanobis-constrained Visual Inversion
While OVI maximizes semantic alignment, it does not guarantee that the resulting embedding $\hat{z}_{x}$ lies on the manifold of realistic image embeddings. To address this, we introduce a regularization term based on the Mahalanobis distance. We pre-compute the mean $\mu_{coco}$ and the inverse covariance matrix $\Sigma^{-1}_{coco}$ of image embeddings from a large reference dataset (MS-COCO). The Mahalanobis loss penalizes embeddings that are far from the center of this distribution:
@@ 4.1 Experimental Setup
Implementation Details. We use the CLIP-ViT-bigG-14 model for text and image encoders and the Kandinsky 2.2 model as our image decoder. All models are frozen during the OVI process. Our primary baseline is ECLIPSE Eclipse, a state-of-the-art resource-efficient prior. We also compare against a strong baseline where the Direct Text Embedding ($z_{y}$) is fed directly to the decoder, which we refer to as TextEmb. We use the MS-COCO (40k images) dataset to compute statistics for our constrained OVI methods.
Evaluation Setting. We evaluate all methods on the T2I-CompBench++ T2ICompBench benchmark, which measures compositional abilities across categories like color, shape, texture, and spatial relations. Unless otherwise specified, all our OVI methods use 1000 optimization steps, 6 pseudo-tokens, and generate 512x512 images with 50 diffusion steps in the decoder. A fixed random seed is used for the decoder to ensure fair comparisons.
@@
  • Attribute Binding (Color, Shape, Texture): Evaluated using a Disentangled BLIP-VQA model, which asks specific questions about object attributes (e.g., “Is the book red?”).
  • 2D Spatial Relations: Evaluated using UniDet to detect objects and verify the geometric relationship between their bounding box centers.
  • 3D Spatial Relations: Evaluated using UniDet combined with a depth estimation model to check for occlusion and relative depth.
  • Non-spatial Relations: Evaluated using CLIPScore, which measures the cosine similarity between the prompt’s text embedding and the generated image’s embedding.
@@ 4.2 Analysis of Unconstrained OVI
This convergence towards the text embedding is quantitatively confirmed by analyzing the cosine similarity, as visualized in Figure 2. The image embedding produced by ECLIPSE consistently maintains a cosine similarity of approximately 0.65 with the text prompt. In contrast, for our unconstrained OVI, the similarity increases logarithmically with the number of optimization steps, eventually plateauing at a value determined by the number of pseudo-tokens. With 1000 steps, we reach a similarity of $\approx$0.6 with 1 token, $\approx$0.84 with 4 tokens, and over 0.9 with 6 tokens. This confirms our hypothesis: the optimization process is effectively pushing the image embedding out of the visual manifold and towards the text manifold, eventually converging to the text embedding itself. This explains why its performance on benchmarks mirrors that of the TextEmb baseline.
@@
Figure 3: Qualitative comparison of OVI variants, the TextEmb baseline, and ECLIPSE. Unconstrained OVI with more tokens produces results visually similar to TextEmb, while ECLIPSE yields higher fidelity.
@@ 4.3 Improving Visual Quality with Priors and Constraints
Figure 4: Evolution of cosine similarity metrics during the optimization process for different OVI configurations. The solid blue curve tracks the similarity between the target text embedding and our optimized OVI embedding. The solid red curve measures the similarity between our OVI embedding and the embedding generated by the ECLIPSE prior. The dashed blue line represents the baseline similarity between the text embedding and the ECLIPSE image embedding. Left: Unconstrained OVI (6 tokens) converges to the text embedding (>0.9>0.9). Center: Mahalanobis constraint ($\lambda_{M}=0.009$) restricts text similarity to $\approx 0.83$. Right: Nearest-Neighbor constraint ($\lambda_{N}=0.5$) stabilizes text similarity around $0.88$ while reaching a similarity of $\approx 0.79$ with ECLIPSE, indicating strong convergence towards the trained prior’s manifold.
@@ 4.3.1 Mahalanobis-constrained OVI
While $\lambda_{M}\approx 0.017$ offered a theoretical balance, we empirically found that a slightly lower value of $\lambda_{M}=0.009$ yielded superior visual results (Figure 7). With this setting, the Mahalanobis loss converges to approximately 28. As visualized in Figure 4 (center), this constraint restricts the text-image cosine similarity to about 0.83, preventing it from reaching the near-perfect alignment (>0.9>0.9) seen in the unconstrained case. This indicates that the optimization is finding a compromise between semantic alignment and the statistical distribution of real images. Despite the visual improvement, the T2I-CompBench++ score dropped to 0.389 (Table 2), further highlighting the trade-off between metric maximization and perceptual quality.
@@ 4.3.2 Nearest-Neighbor Constrained OVI
Figure 6: Nearest neighbors retrieved from the MS-COCO dataset for six different text prompts. These real images serve as the reference “anchors” ($z_{closest}$) for our Nearest-Neighbor Constrained OVI, guiding the optimization towards realistic visual compositions found in the dataset.
Figure 7: Qualitative comparison of constrained OVI methods. Both Mahalanobis and Nearest-Neighbor (NN) constraints improve visual quality over the base OVI and TextEmb baselines. The NN-constrained OVI produces images that are visually competitive with the trained ECLIPSE prior.
@@ 5 Discussion and Limitations
The Metric Problem: High Scores vs. Visual Quality. Our most striking finding is the discrepancy between the quantitative scores from T2I-CompBench++ and the perceived visual quality of the generated images. The TextEmb baseline, despite its lower visual fidelity, consistently outperforms a strong trained prior like ECLIPSE. This is not because the metrics are simply “coupled with raw CLIP similarity,” but due to a more nuanced reason revealed by our analysis.
The benchmark metrics (such as those based on Disentangled BLIP-VQA) often verify the presence of atomic attributes (e.g., “is the book red?” or “is there a car?”) and the raw text embedding is the purest representation of these attributes, hence gives the highest score. Instead, a prior projects this representation into the natural image manifold. In this sense, the text embedding is well-suited for satisfying discrete directives, while the value of a true prior (trained, like ECLIPSE, or optimized, like our constrained OVI) lies in integrating these factors into a unified and perceptually rich composition. This highlights the need for holistic evaluation metrics that balance atomic correctness with overall visual quality.
Limitations. Although promising, our OVI approach has limitations. Its primary drawback is inference speed. The iterative optimization process (1000 steps) is orders of magnitude slower than a single forward pass through a trained prior like ECLIPSE. Secondly, our constrained methods depend on an external dataset (MS-COCO). While deriving statistics for the Mahalanobis constraint is a one-time pre-computation, the Nearest-Neighbor constraint introduces additional computational overhead during inference. Specifically, for every new prompt, the method requires to compute the cosine similarity between the input text embedding and the entire reference dataset to identify the closest match.

### v3

@@ Abstract
Diffusion models have established the state-of-the-art in text-to-image generation, but their performance often relies on a diffusion prior network to translate text embeddings into the visual manifold for easier decoding. These priors are computationally expensive and require extensive training on massive datasets. In this work, we challenge the necessity of a trained prior at all by employing Optimization-based Visual Inversion (OVI), a training-free and zero-shot alternative, to replace the need for a prior. OVI initializes a latent visual representation from random pseudo-tokens and iteratively optimizes it to maximize the cosine similarity with the input textual prompt embedding. We further propose two novel constraints, a Mahalanobis-based and a Nearest-Neighbor loss, to regularize the OVI optimization process toward the distribution of realistic images. Our experiments, conducted on Kandinsky 2.2, show that OVI can serve as an alternative to traditional priors. More importantly, our analysis reveals a critical flaw in current evaluation benchmarks like T2I-CompBench++, where simply using the text embedding as a prior achieves surprisingly high scores, despite lower perceptual quality. Our constrained OVI methods improve visual fidelity over this baseline, with the Nearest-Neighbor approach proving particularly effective. It achieves quantitative scores comparable to or higher than the state-of-the-art data-efficient prior, underscoring the potential of optimization-based strategies as viable, training-free alternatives to traditional priors. The code will be publicly available upon acceptance.
@@ 1 Introduction
Text-to-image (T2I) generation has seen remarkable progress, evolving from autoregressive approaches like DALL-E ramesh2021zero to diffusion models ho2020denoising; StableDiffusion, which now represent the state-of-the-art. Architectures like unCLIP unCLIP have demonstrated that a hierarchical approach, which first generates a CLIP image embedding radford2021learning from a text caption using a prior model and then generates an image from that embedding using a decoder, can produce high-fidelity and diverse images. However, this performance comes at a significant cost: the diffusion prior is typically a large model that is expensive to train, requiring billions of image-text pairs and substantial computational resources.
@@
Recent work has focused on making these priors more efficient. ECLIPSE Eclipse, for example, introduced a data-efficient prior that achieves competitive results at a fraction of the parameters and training samples. Despite these advancements, the fundamental paradigm remains unchanged: the prior must still be learned through an extensive training phase, limiting accessibility and adaptability.
@@
In this work, we investigate a radical alternative: “can we entirely eliminate the training phase for the prior?” Building on recent work on modality inversion CrossTheGap, we leverage Optimization-based Visual Inversion (OVI), a training-free method proposed therein to cast intra-modal tasks into inter-modal ones, to substitute the prior with an on-the-fly optimization process. Instead of relying on a mapping network, OVI generates a CLIP image embedding by iteratively refining a random latent vector to maximize its semantic alignment with the target textual prompt embedding. To further guide this process towards plausible visual representations, we introduce two novel regularization strategies: a Mahalanobis constraint lee2018simple that pulls the embedding towards the global distribution of real image embeddings, and a Nearest-Neighbor constraint, inspired by retrieval-augmented approaches blattmann2022retrieval; chen2022re; sheynin2022knn, that aligns it with the most similar real image embedding from a reference dataset.
@@
Our preliminary results reveal several interesting aspects of T2I generation. We find that unconstrained OVI produces qualitatively similar results to directly using the text embedding in the decoder, a surprisingly strong baseline first noted in unCLIP. This baseline achieves state-of-the-art scores on the benchmark T2I-CompBench++ T2ICompBench, often exceeding expensive trained priors like ECLIPSE. This suggests a potential “metric problem,” consistent with the misalignment of metrics like FID heusel2017gans and CLIPScore hessel2021clipscore with human preference pickscore. Specifically, current benchmarks seem to over-reward semantic similarity at the expense of visual realism. Our constrained OVI methods improve visual fidelity over this baseline, with the Nearest-Neighbor approach proving particularly effective. It achieves quantitative scores comparable to or higher than the state-of-the-art prior ECLIPSE while utilizing a reference set of only 40k images ($\approx$0.8% of the 5 million used by the latter), highlighting the potential of optimization-based strategies as resource-efficient, training-free alternatives to data-hungry learned priors.
@@ 2.1 Diffusion Models for Text-to-Image Generation
Diffusion models have become the dominant architecture for T2I synthesis. The unCLIP unCLIP framework popularized a two-stage pipeline consisting of a prior and a decoder. The decoder, often a U-Net based model, is trained to generate an image conditioned on a CLIP image embedding. Our work leverages the publicly available Kandinsky 2.2 decoder Kandinsky, which is known for its high-quality image generation capabilities. While recent adapters like ControlNet zhang2023adding enable fine-grained structural control, our work focuses on the semantic guidance provided by the global image prior.
@@ 2.2 Image Priors for Text-to-Image Generation
The prior model is responsible for mapping the textual prompt embedding $z_{y}$ to the corresponding image embedding $z_{x}$. The original unCLIP unCLIP explored both autoregressive and diffusion-based priors, with the latter generally proving more efficient and effective. More recently, ECLIPSE Eclipse proposed a lightweight, non-diffusion prior trained with a combination of a projection and a contrastive loss, significantly reducing computational requirements while maintaining high performance. Our work investigates the possibility of omitting the prior training phase, proposing a zero-shot, optimization-based substitute.
@@ 2.3 Modality Inversion and Optimization-based Methods
The idea of inverting a model or optimizing in a latent space is not new. The concept of test-time optimization has roots in the seminal Deep Image Prior ulyanov2018deep, while more recently, Null-text Inversion mokady2023null applied pivotal tuning to the unconditional embedding for editing real images. In the context of CLIP, CrossTheGap investigated modality inversion by adapting Textual Inversion (OTI) gal2022image and introducing Optimization-based Visual Inversion (OVI) to expose and bridge the gap between the text and image embedding manifolds liang2022mind. In parallel, methods like SEARLE SEARLE have used optimization to project a reference image into a “pseudo-token” for composed image retrieval, demonstrating the power of training-free optimization to achieve strong multimodal alignment. Our OVI-based approach builds on these concepts, adapting modality inversion to serve as a generative, training-free prior for text-to-image generation.
@@ 3.2 Optimization-based Visual Inversion (OVI)
The optimization-based approach frames the prior’s task as an optimization problem. Instead of learning a mapping, we directly search for an image embedding $\hat{z}_{x}$ that aligns with the given text embedding $z_{y}$. We initialize a set of $N$ learnable pseudo-tokens, which form our latent image embedding $z_{img}^{(0)}$. We then iteratively update this embedding for $T$ steps using an optimizer to minimize a loss function based on cosine similarity with the target text embedding, as computed by the CLIP encoders:
@@ 3.3 Mahalanobis-constrained Visual Inversion
While OVI maximizes semantic alignment, it does not guarantee that the resulting embedding $\hat{z}_{x}$ lies on the manifold of realistic image embeddings. To address this, we introduce a regularization term based on the Mahalanobis distance. We pre-compute the mean $\mu_{coco}$ and the inverse covariance matrix $\Sigma^{-1}_{coco}$ of image embeddings from a large reference dataset (MS-COCO lin2014microsoft). The Mahalanobis loss penalizes embeddings that are far from the center of this distribution:
@@ 4.1 Experimental Setup
Implementation Details. We use the OpenCLIP cherti2023reproducible implementation of the ViT-bigG-14 model (trained on LAION-5B schuhmann2022laion) for text and image encoders and the Kandinsky 2.2 Kandinsky model as our image decoder. All models are frozen during the OVI process. Our primary baseline is ECLIPSE Eclipse, a state-of-the-art resource-efficient prior. We also compare against a strong baseline where the Direct Text Embedding ($z_{y}$) is fed directly to the decoder, which we refer to as TextEmb. We use the MS-COCO (40k images) dataset lin2014microsoft to compute statistics for our constrained OVI methods.
Evaluation Setting. We evaluate all methods on the T2I-CompBench++ T2ICompBench benchmark, which measures compositional abilities across categories like color, shape, texture, and spatial relations. Unless otherwise specified, all our OVI methods use 1000 optimization steps, 6 pseudo-tokens, and generate 512x512 images with 50 diffusion steps (DDIM song2021denoising) using Classifier-Free Guidance ho2022classifier with a scale of 4.0 in the decoder. A fixed random seed is used for the decoder to ensure fair comparisons.
@@
Zero tensor
ECLIPSE pipeline’s default method
@@
  • Attribute Binding (Color, Shape, Texture): Evaluated using a Disentangled BLIP-VQA model li2022blip, which asks specific questions about object attributes (e.g., “Is the book red?”).
  • 2D Spatial Relations: Evaluated using UniDet zhou2022simple to detect objects and verify the geometric relationship between their bounding box centers.
  • 3D Spatial Relations: Evaluated using UniDet combined with the DPT ranftl2021vision depth estimation model to check for occlusion and relative depth.
  • Non-spatial Relations: Evaluated using CLIPScore hessel2021clipscore, which measures the cosine similarity between the prompt’s text embedding and the generated image’s embedding.
@@ 4.2 Analysis of Unconstrained OVI
Figure 2: Qualitative comparison of OVI variants, the TextEmb baseline, and ECLIPSE Eclipse. Unconstrained OVI with more tokens produces results visually similar to TextEmb, while ECLIPSE yields higher fidelity.
This convergence towards the text embedding is quantitatively confirmed by analyzing the cosine similarity, as visualized in Figure 3. The image embedding produced by ECLIPSE consistently maintains a cosine similarity of approximately 0.65 with the text prompt. In contrast, for our unconstrained OVI, the similarity increases logarithmically with the number of optimization steps, eventually plateauing at a value determined by the number of pseudo-tokens. With 1000 steps, we reach a similarity of $\approx$0.63 with 1 token, $\approx$0.84 with 3 tokens, $\approx$0.93 with 6 tokens, and $\approx$0.96 with 9 tokens. This confirms our hypothesis: the optimization process is effectively pushing the image embedding out of the visual manifold and towards the text manifold, eventually converging to the text embedding itself. This explains why its performance on benchmarks mirrors that of the TextEmb baseline.
1 TokenRefer to caption
3 TokensRefer to caption
6 TokensRefer to caption
9 TokensRefer to caption
@@ 4.3.1 Mahalanobis-constrained OVI
$\lambda=0.05$
$\lambda=0.1$
a plastic toy and a glass bottle
an oblong cucumber and a teardrop plum
@@
While $\lambda_{M}\approx 0.017$ offered a theoretical balance, we empirically found that a slightly lower value of $\lambda_{M}=0.009$ yielded superior visual results (Figure 5). With this setting, the Mahalanobis loss converges to approximately 28. As visualized in Figure 6 (center), this constraint restricts the text-image cosine similarity to about 0.80, preventing it from reaching the near-perfect alignment (>0.9>0.9) seen in the unconstrained case. This indicates that the optimization is finding a compromise between semantic alignment and the statistical distribution of real images. Despite the visual improvement, the T2I-CompBench++ score dropped to 0.389 (Table 2), further highlighting the trade-off between metric maximization and perceptual quality.
Figure 5: Qualitative comparison of constrained OVI methods. Both Mahalanobis and Nearest-Neighbor ($N$) constraints improve visual quality over the base OVI and TextEmb baselines. The NN-constrained OVI produces images that are visually competitive with the trained ECLIPSE Eclipse prior.
Ovi (6 tokens)
Ovi + Mahalanobis ($\lambda=0.009$)
Ovi + N-N ($\lambda=0.5$)
Figure 6: Evolution of cosine similarity metrics during the optimization process for different OVI configurations. The solid blue curve tracks the similarity between the target text embedding and our optimized OVI embedding. The solid red curve measures the similarity between our OVI embedding and the embedding generated by the ECLIPSE Eclipse prior. The dashed blue line represents the baseline similarity between the text embedding and the ECLIPSE image embedding. Top: Unconstrained OVI (6 tokens) converges to the text embedding (>0.9>0.9). Bottom-Left: Mahalanobis constraint ($\lambda_{M}=0.009$) restricts text similarity to $\approx 0.80$. Bottom-Right: Nearest-Neighbor constraint ($\lambda_{N}=0.5$) stabilizes text similarity around $0.88$ while reaching a similarity of $\approx 0.79$ with ECLIPSE, indicating strong convergence towards the trained prior’s manifold.
@@ 4.3.2 Nearest-Neighbor Constrained OVI
a green bench and a blue bowl
a giraffe next to a lamp
a plastic toy and a glass bottle
a dog in front of a desk
an oblong cucumber and a teardrop plum
a person is looking at a rainbow and        marveling at its beauty
Figure 7: Nearest neighbors retrieved from the MS-COCO lin2014microsoft dataset for six different text prompts. These real images serve as the reference “anchors” ($z_{closest}$) for our Nearest-Neighbor Constrained OVI, guiding the optimization towards realistic visual compositions found in the dataset.
@@ 5 Discussion and Limitations
The Metric Problem: High Scores vs. Visual Quality. Our most striking finding is the discrepancy between the quantitative scores from T2I-CompBench++ and the perceived visual quality of the generated images. The TextEmb baseline, despite its lower visual fidelity, consistently outperforms a strong trained prior like ECLIPSE Eclipse. This is not because the metrics are simply “coupled with raw CLIP similarity,” but due to a more nuanced reason revealed by our analysis.
The benchmark metrics (such as those based on Disentangled BLIP-VQA li2022blip) often verify the presence of atomic attributes (e.g., “is the book red?” or “is there a car?”) and the raw text embedding is the purest representation of these attributes, hence gives the highest score. Instead, a prior projects this representation into the natural image manifold. In this sense, the text embedding is well-suited for satisfying discrete directives, while the value of a true prior (trained, like ECLIPSE, or optimized, like our constrained OVI) lies in integrating these factors into a unified and perceptually rich composition. This highlights the need for holistic evaluation metrics that balance atomic correctness with overall visual quality.
Limitations. Although promising, our OVI approach has limitations. Its primary drawback is inference speed. The iterative optimization process (1000 steps) is orders of magnitude slower than a single forward pass through a trained prior like ECLIPSE. Secondly, our constrained methods depend on an external dataset (MS-COCO). While deriving statistics for the Mahalanobis constraint is a one-time pre-computation, the Nearest-Neighbor constraint introduces additional computational overhead during inference. Specifically, for every new prompt, the method requires computing the cosine similarity between the input text embedding and the entire reference dataset to identify the closest match.